A numerical algorithm approximates the value of sin(0.5) as 0.465. Note that is 0.5 radians. The true relative error (to 4 decimal places) in this approximation is
Answer #1

Sin(0.5 radians) = sin(28.6478898∘)≈0.479425539

Relative error is given as:

(measured value - true value) / (true value)

= (0.465 - 0.479425539) / 0.479425539 = -0.0301